OpenStudy (anonymous):

write the word or phrase that best completes each statement or answers the question. The radius of the Earth is R. At what distance above the Earth's surface will the acceleration of gravity be at 4.9 m/s^2

OpenStudy (anonymous):

How can you answer this with a word or phrase?

OpenStudy (anonymous):

i believe the answer would be 9.02 * 10^6 is that correct?

OpenStudy (anonymous):

or i guess it would be (9.02 * 10^6)-R

OpenStudy (ivancsc1996):

The force of gravity is \[F _{g}=\frac{ m m _{earth} G }{ r ^{2} }=m g \rightarrow g=\frac{ m _{earth} G }{ r ^{2} } \rightarrow r=\sqrt{\frac{ m _{earth}G }{ g }}\] were g is 4.9 meters per second square, m earth is the mass of the earth, G is the gravitational constant and r is the height

OpenStudy (anonymous):

thanks then i think I got it.